• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 13:26:33 on June12,2005

    Entry number 145939

    This entry is a followup to: 145938

    keyword=coinc. timing is ok

    All I did (halog 145555) was move the randoms over. I didn't include
    the plot that proves all is well because I figured anyone could see
    it was an improvement. But anyway, here it is (fig 1) for run 4426.
    Compared with the D.ctimel spectrum before the adjustment (halog
    145508) you can see why I made the change and that it was indeed
    an improvement.

    Note:
    a) The physical peaks did not move -- as expected. They go where
    they go due to fixed cable lengths.
    b) The width of the randoms is the same -- as expected. That is
    determined by the widths of T1 and T3 which I did not change.
    c) The randoms moved over 16 nsec to the right -- as expected. This
    gives more "breathing room" for the peaks. Admittedly, if your
    physics was at channel <= 1500 in this distribution then it did get
    chopped off by the change. But I suppose that's not your problem.

    Fig 1 is for run 4426.


    FIGURE 1